# Non-Clifford and parallelizable fault-tolerant logical gates on constant and almost-constant rate homological quantum LDPC codes via higher symmetries

### v2

@@ II.2 Logical non-Clifford gates and the triple-intersection invariant
In order to label the logical qubits, we choose a particular 2nd homology basis $B_{2}=\{\alpha_{2}\}$. Therefore, we can label the logical qubits with the label of the logical-$X$ membranes, i.e., $(\alpha_{2};i)$ with $\alpha_{2}\in B_{2}$. We note that the dual logical-$Z$ string of the logical-$X$ membrane $\overline{X_{\alpha_{2};i}}$ ($\overline{X}_{\alpha_{2}}^{(i)}$) is $\overline{Z_{\alpha_{1};i}}$ ($\overline{Z}_{\alpha_{1}}^{(i)}$) satisfying $|\alpha_{1}\cap\alpha_{2}|=1$ and $|\alpha_{1}\cap\alpha^{(2)}_{2}|=0$ for any $\alpha^{(2)}_{2}\in B_{2}$ such that $\alpha^{(2)}_{2}\neq\alpha_{2}$, where ‘$|\cdot\cap\cdot|\in\mathbb{Z}_{2}$’ represents the $\mathbb{Z}_{2}$ intersection number, i.e., the number of intersection points mod 2  [^3] A more general quantity to be considered is the algebraic intersection number $|\alpha_{1}\cap\beta_{2}|_{\mathbb{Z}}\in\mathbb{Z}$, which has a ‘$\pm$’ sign for each intersection point depending the orientation of the cycles and is a topological invariant. In contrast, geometric intersection number only associate a ‘$+$’ sign for each intersection point, and is hence not topological but depending the choice of the representative of each cycle. In the $\mathbb{Z}_{2}$ case, $+^{\prime}$ and $-^{\prime}$ signs are identified and it turns out the algebraic and geometric intersection numbers mod 2 are the same. For more general case like the $\mathbb{Z}_{N}$ toric code, one should consider algebraic intersection number mod $N$.. We see that $\alpha_{1}$ is just the Poincaré dual cycle of $\alpha_{2}$  [^4] Strictly speaking Poincaré dual is between the cycle $\alpha_{2}$ and the cocycle $\alpha^{1}$. In this paper, we sometimes relax the use of this terminology by calling $\alpha^{1}$ the Poincaré dual cycle of $\alpha_{2}$.. From now on, we use the same Greek letter to denote the pair of Poincaré dual cycles. The dual pairs obey the anti-commutation relations $\overline{X_{\alpha_{2};i}}\ \overline{Z_{\alpha_{1};i}}$$=$$-\overline{Z_{\alpha_{1};i}}\ \overline{X_{\alpha_{2};i}}$ and $\overline{X}_{\alpha_{2}}^{(i)}\overline{Z}_{\alpha_{1}}^{(i)}=-\overline{Z}_{\alpha_{1}}^{(i)}\overline{X}_{\alpha_{2}}^{(i)}$. We hence also obtain the corresponding 1st homology basis $B_{1}=\{\alpha_{1}\}$. Therefore, both labels $(\alpha_{2};i)$ and $(\alpha_{1};i)$ correspond to the same logical qubit.
@@ II.3 Parallelizable logical Clifford gates
From the discussion in Fig. 3, we can see that transversal CZ gate can be applied only on a particular codimension-1 (2-dimensional) submanifold, i.e., a particular membrane (2-cycle) $\alpha_{2}$.
@@ II.4 Interaction hypergraph and hypergraph magic state injection
We start with a 3-manifold $\mathcal{M}^{3}$ and with particular choice of the 2nd homology basis $B_{2}$. Its $\mathbb{Z}_{2}$ triple intersection number $|\alpha_{2}^{(i)}\cap\alpha_{2}^{(j)}\cap\alpha_{2}^{(k)}|\in\mathbb{Z}_{2}$ with $\alpha_{2}^{(i)},\alpha_{2}^{(j)},\alpha_{2}^{(k)}\in B_{2}$ can be encoded into a base interaction hypergraph, or synonymously an intersection hypergraph, as shown in Fig. 4(a). The hypergraph $G_{h}$$=$$(V,E_{h})$ is composed of a set of vertices $V$ (represented by grey circles) and a set of hyperedge $E_{h}$ (represented by a tri-junction with three edges meeting at a square node) connecting triplets of vertices. Each vertex $v\in V$ corresponds to a 2-cycle $\alpha_{2}^{(i)}\in B_{2}$. Each hyperedge $e_{h}\in E_{h}$ corresponds to a $\mathbb{Z}_{2}$ triple intersection point, which is connected to three vertices with their represented 2-cycles satisfying the condition $|\alpha_{2}^{(i)}\cap\alpha_{2}^{(j)}\cap\alpha_{2}^{(k)}|=1$. Note that as illustrated in Fig. 4(a), the base interaction hypergraph does not have to be a planar hypergraph since the hyperedges can have crossings.
@@
Quite interestingly, the full-interaction hypergraph has a one-to-one correspondence with the quantum hypergraph states [50], which can be considered as the generalization of graph states. One can define a $q$-uniform hypergraph state on a hypergraph $G_{h}=(V,E_{h})$ as follows. One assigns to each vertex a qubit initialized in the $\ket{+}$ state. For each $q$-hyperedge, one perform a $\text{C}^{q}\text{Z}$ gate between the $q$ connected qubits (vertices) labeled by $v_{1},v_{2},...,v_{q}$. We hence reach the following $q$-uniform hypergraph state:
|$\ket{g_{q}}=\prod_{\{v_{1},v_{2},...,v_{q}\}\in E_{h}}\text{C}^{q}\text{Z}(v_{1},v_{2},...,v_{q})\ket{+}^{\otimes m},$ (17)
@@
Figure 5: Correspondence between an interaction hypergraph of a color code on a 3-manifold in (a) with a logical 3-uniform hypergraph magic state in (b) injected via a transversal $T$ gate acted on the entire code with each logical qubit initialized in state $\overline{\ket{+}}$. The number of logical CCZ gates in the quantum circuit in (b) equaling the number hyperedges in the corresponding hypergraph shown in (a) defines the complexity of the hypergraph state.
In the case of the color code defined on 3-manifold, we have $q=3$. If we initialize each logical qubit in the logical state $\overline{\ket{+}}$ and then apply the transversal $T$ gate, we arrive the following 3-uniform logical hypergraph state according to Eq. (14):
|$\overline{\ket{g_{3}}}=$ $\prod_{\alpha_{2},\beta_{2},\gamma_{2}\in B_{2}}[\overline{\text{CCZ}}((\alpha_{2};1),(\beta_{2};2),(\gamma_{2};3))]^{|\alpha_{2}\cap\beta_{2}\cap\gamma_{2}|}$
|+¯kabsentsuperscript¯$\qquad\qquad\cdot\overline{\ket{+}}^{\otimes k}$
$=$ $\prod_{\{v_{1},v_{2},v_{3}\}\in E_{h}}\overline{\text{CCZ}}(v_{1},v_{2},v_{3})\overline{\ket{+}}^{\otimes k}.$ (18)
@@
Note that the hypergraph state contains so-called ‘magic[51], since it is a non-stabilizer state beyond the description of the stabilizer formalism. In our context, the logical hypergraph state in Eq. (II.4) is a non-stabilizer state at the logical level, even though the microscopic wavefunction is in the code space a stabilizer code. Therefore, this hypergraph state is a specific type of magic state which is a valuable computational resource for the magic state injection and distillation protocols [53]. We note that instead of slowly doing extensive rounds of magic state distillation starting from a low-fidelity magic state at $O(1)$ distance, here we can directly inject a high-fidelity hypergraph magic state at $O(d)$ distance. The injected hypergraph magic state can either be used as resources for gate-based or measurement-based quantum computation schemes [52] at the logical level.
@@
III Logical gates for three identical copies of toric codes on 3D manifold via constant-depth circuits and the connection to emergent $k$-form symmetries in topological quantum field theory
In this section, we derive the logical gate expressions Eq. (14) and (16) in a more fundamental way using an operator-valued cochain formalism originated from (3+1)D topological quantum field theory (TQFT). This derivation is independent of the detailed microscopic lattice models one chooses such as a color code or multiple copies of toric codes, or any lattice variation of these models. In particular, the logical gate operators originate from $k$-form emergent symmetries expressed as TQFT path integrals and mathematically correspond to cohomology operation via cup products. In this way, the triple intersection numbers naturally arise as topological invariants in the TQFT path integral.
Even more interestingly, this operator-valued cochain formalism naturally adapts to a discrete lattice or trianulation. This allows us to explicitly construct constant-depth circuits acting on three identical copies of toric codes defined on an arbitrary triangulation of arbitrary closed 3-manifolds. This is different form the transversal CCZ [62] acted on three different toric codes, with two of them having larger stabilizer weights than the standard cubic lattice 3D toric-code. Our construction hence allows implementing logical CCZ gate with lower-weight stabilizer. More detailed study of this formalism including the logical gates on arbitrary lattices such as cubical lattices and its application to a broader class of cohomology operation will be presented in an upcoming paper [48].
@@ III.1 Operator-valued cochain formalism
[…]
@@
Note that the Pauli operator has eigenvalues ±$\pm 1$ instead. Physically, $a^{(i)}$ corresponds to the $\mathbb{Z}_{2}$ electric gauge field. Similarly, there also exists a 1-cochain $b^{(i)}$ corresponding to the magnetic $\mathbb{Z}_{2}$ gauge field, which corresponds to a Pauli-X operator on each edge $e$ as
@@ III.2 Logical CCZ as a 0-form symmetry
For a triangulated manifold, one can define the cup product $\cup$’ between a $p$-cochain $\alpha^{p}$ and $q$-cochain $\beta^{q}$ evaluated on a $(p+q)$-simplex $[v_{0}v_{1}\cdots v_{p+q}]$ as [63]
@@
The cup-product evaluated on each 3-simplex $[v_{0}v_{1}v_{2}v_{3}]$ ($v_{i}$ represents the four vertices in the simplex)
@@
which explicitly shows the corresponding constant-depth circuit composed of many CCZ gates coupling the qubits in three copies of toric codes located on edge $[v_{0}v_{1}],[v_{1}v_{2}]$ and $[v_{2}v_{3}]$ in each simplex $[v_{0}v_{1}v_{2}v_{3}]$, as illustrated in Fig. 7.
Figure 7: Illustration of the evaluation of the cup product in Eq. (III.2). A CCZ gate is coupling the qubits located on red, blue and green edges, i.e., $[v_{0}v_{1}],[v_{1}v_{2}],[v_{2}v_{3}]$ in each simplex labeled by $[v_{0}v_{1}v_{2}v_{3}]$. The CCZ gates in neighboring simplices may have overlap and they hence compose a constant-depth local circuit instead of a strictly transversal gate.
@@
We consider an $X$-stabilizer $S^{(1)}_{X;v}$$=$$\prod_{e\supset v}X^{(1)}(e)$ located at vertex $v$ in the 1st copy of the toric code. We then consider conjugating $U$ with $S^{(1)}_{X;v}$:
@@
The condition in Eq. (37) should be considered as the gauge-invariance condition in the code space, since $a^{(1)}\rightarrow a^{(1)}+d\tilde{v}$ is just a gauge transformation in the $\mathbb{Z}_{2}$ gauge theory  [^6] This is in analogy to the famous gauge transformation $\vec{a}\rightarrow\vec{a}+\nabla\chi$ in Maxwell’s theory or the corresponding gauge theory in the continuum, where $\vec{a}$ is the quantum vector potential. The gauge invariance of the magnetic field $B=\nabla\times\vec{a}=0$ is satisfied due to the Stokes theorem: ×$\nabla\times\nabla\chi=0$.
@@
[…]
@@
Mathematically, the gauge invariance condition is associated with topological invariance of $U$ with respect to arbitrary deformation of the cocycles by coboundaries.
@@
$a^{(1)}=\hat{n}\alpha^{1},\quad a^{(2)}=\hat{m}\beta^{1},\quad a^{(3)}=\hat{l}\gamma^{1},$ (41)
where the quantum variable $\hat{n}$, $\hat{m}$ and $\hat{l}$ with eigenvalues $\{0,1\}$ are the winding numbers, i.e., coefficients of the cocycle basis. We can hence re-express $U$ as
$U=$ $(-1)^{\int_{\mathcal{L}}(\hat{n}\alpha^{1})\cup(\hat{m}\beta^{1})\cup(\hat{l}\gamma^{1})}=\left[(-1)^{\hat{n}\hat{m}\hat{l}}\right]^{\int_{\mathcal{L}}\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}$
$=$ $\overline{\text{CCZ}}[(\alpha^{1};1),(\beta^{1};2),(\gamma^{1};3)]^{\int_{\mathcal{L}}{\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}}$
$=$ $\overline{\text{CCZ}}[(\alpha_{2};1),(\beta_{2};2),(\gamma_{2};3)]^{{|\alpha_{2}\cup\beta_{2}\cup\gamma_{2}|}},$ (42)
where the third equality has use the relation $(-1)^{\hat{n}\hat{m}\hat{l}}$$=$$\overline{\text{CCZ}}$. Here, $(\alpha^{1};1)$, $(\beta^{1};2)$ and $(\gamma^{1};3)$ are the labels of the logical qubit using the cocycle basis and the copy number. The logical CCZ gate between logical qubits $(\alpha^{1};1)$, $(\beta^{1};2)$ and $(\gamma^{1};3)$ is only non-trivial (not logical identity) if and only if $\int_{\mathcal{L}}{\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}$ evaluates non-trivially, i.e., have a non-trivial triple intersection of the Poincare dual: $|\alpha_{2}\cup\beta_{2}\cup\gamma_{2}|=1$.

### v3

@@ II.2 Logical non-Clifford gates and the triple-intersection invariant
In order to label the logical qubits, we choose a particular 2nd homology basis $B_{2}=\{\alpha_{2}\}$. Therefore, we can label the logical qubits with the label of the logical-$X$ membranes, i.e., $(\alpha_{2};i)$ with $\alpha_{2}\in B_{2}$. We note that the dual logical-$Z$ string of the logical-$X$ membrane $\overline{X_{\alpha_{2};i}}$ ($\overline{X}_{\alpha_{2}}^{(i)}$) is $\overline{Z_{\alpha_{1};i}}$ ($\overline{Z}_{\alpha_{1}}^{(i)}$) satisfying $|\alpha_{1}\cap\alpha_{2}|=1$ and $|\alpha_{1}\cap\alpha^{(2)}_{2}|=0$ for any $\alpha^{(2)}_{2}\in B_{2}$ such that $\alpha^{(2)}_{2}\neq\alpha_{2}$, where $|\cdot\cap\cdot|\in\mathbb{Z}_{2}$’ represents the $\mathbb{Z}_{2}$ intersection number, i.e., the number of intersection points mod 2  [^3] A more general quantity to be considered is the algebraic intersection number $|\alpha_{1}\cap\beta_{2}|_{\mathbb{Z}}\in\mathbb{Z}$, which has a $\pm$’ sign for each intersection point depending the orientation of the cycles and is a topological invariant. In contrast, geometric intersection number only associate a ‘++’ sign for each intersection point, and is hence not topological but depending the choice of the representative of each cycle. In the $\mathbb{Z}_{2}$ case, $^{\prime}$ and $^{\prime}$ signs are identified and it turns out the algebraic and geometric intersection numbers mod 2 are the same. For more general case like the $\mathbb{Z}_{N}$ toric code, one should consider algebraic intersection number mod $N$.. We see that $\alpha_{1}$ is just the Poincaré dual cycle of $\alpha_{2}$  [^4] Strictly speaking Poincaré dual is between the cycle $\alpha_{2}$ and the cocycle $\alpha^{1}$. In this paper, we sometimes relax the use of this terminology by calling $\alpha^{1}$ the Poincaré dual cycle of $\alpha_{2}$.. From now on, we use the same Greek letter to denote the pair of Poincaré dual cycles. The dual pairs obey the anti-commutation relations $\overline{X_{\alpha_{2};i}}\ \overline{Z_{\alpha_{1};i}}$$\overline{Z_{\alpha_{1};i}}\ \overline{X_{\alpha_{2};i}}$ and $\overline{X}_{\alpha_{2}}^{(i)}\overline{Z}_{\alpha_{1}}^{(i)}=-\overline{Z}_{\alpha_{1}}^{(i)}\overline{X}_{\alpha_{2}}^{(i)}$. We hence also obtain the corresponding 1st homology basis $B_{1}=\{\alpha_{1}\}$. Therefore, both labels $(\alpha_{2};i)$ and $(\alpha_{1};i)$ correspond to the same logical qubit.
@@ II.3 Parallelizable logical Clifford gates
From the discussion in Sec. II.2 , we can see that transversal CZ gate can be applied only on a particular codimension-1 (2-dimensional) submanifold, i.e., a particular membrane (2-cycle) $\alpha_{2}$, as shown in Fig. 3.
@@ II.4 Interaction hypergraph and hypergraph magic state injection
We start with a 3-manifold $\mathcal{M}^{3}$ and with particular choice of the 2nd homology basis $B_{2}$. Its $\mathbb{Z}_{2}$ triple intersection number $|\alpha_{2}^{(i)}\cap\alpha_{2}^{(j)}\cap\alpha_{2}^{(k)}|\in\mathbb{Z}_{2}$ with $\alpha_{2}^{(i)},\alpha_{2}^{(j)},\alpha_{2}^{(k)}\in B_{2}$ can be encoded into a base interaction hypergraph, or synonymously an intersection hypergraph, as shown in Fig. 4(a). The hypergraph $G_{h}$$(V,E_{h})$ is composed of a set of vertices $V$ (represented by grey circles) and a set of hyperedge $E_{h}$ (represented by a tri-junction with three edges meeting at a square node) connecting triplets of vertices. Each vertex $v\in V$ corresponds to a 2-cycle $\alpha_{2}^{(i)}\in B_{2}$. Each hyperedge $e_{h}\in E_{h}$ corresponds to a $\mathbb{Z}_{2}$ triple intersection point, which is connected to three vertices with their represented 2-cycles satisfying the condition $|\alpha_{2}^{(i)}\cap\alpha_{2}^{(j)}\cap\alpha_{2}^{(k)}|=1$. Note that as illustrated in Fig. 4(a), the base interaction hypergraph does not have to be a planar hypergraph since the hyperedges can have crossings.
@@
[…]
@@
[…]
@@
Note that the hypergraph state contains so-called ‘magic[55], since it is a non-stabilizer state beyond the description of the stabilizer formalism. In our context, the logical hypergraph state in Eq. (II.4) is a non-stabilizer state at the logical level, even though the microscopic wavefunction is in the code space a stabilizer code. Therefore, this hypergraph state is a specific type of magic state which is a valuable computational resource for the magic state injection and distillation protocols [57]. We note that instead of slowly doing extensive rounds of magic state distillation starting from a low-fidelity magic state at $O(1)$ distance, here we can directly inject a high-fidelity hypergraph magic state at $O(d)$ distance without distillation. We hence call this scheme a ‘magic state fountain’. The injected hypergraph magic state can either be used as resources for gate-based or measurement-based quantum computation schemes [56] at the logical level.
@@
III Logical gates for three identical copies of toric codes on 3D manifold via constant-depth circuits and the connection to emergent $k$-form symmetries in TQFT
In this section, we derive the logical gate expressions Eq. (14) and (16) in a more fundamental way using an operator-valued cochain formalism originated from (3+1)D topological quantum field theory (TQFT) (see also Ref. [66] for the application of this formalism). This derivation is independent of the detailed microscopic lattice models one chooses such as a color code or multiple copies of toric codes, or any lattice variation of these models. In particular, the logical gate operators originate from $k$-form emergent symmetries expressed as TQFT path integrals and mathematically correspond to cohomology operation via cup products. In this way, the triple intersection numbers naturally arise as topological invariants in the TQFT path integral.
Even more interestingly, this operator-valued cochain formalism naturally adapts to a discrete lattice or trianulation. This allows us to explicitly construct constant-depth circuits acting on three identical copies of toric codes defined on an arbitrary triangulation of arbitrary closed 3-manifolds. This is different form the transversal CCZ [47] acted on three different toric codes, with two of them having larger stabilizer weights than the standard cubic lattice 3D toric-code. Our construction hence allows implementing logical CCZ gate with lower-weight stabilizer. More elaborated study of this formalism including the logical gates on arbitrary lattices such as cubical lattices and its application to a broader class of cohomology operation will be presented in the follow-up papers [51, 52].
@@ III.1 Operator-valued cochain formalism
We now introduce an operator-valued cochain formalism to describe the operators in the code, which originates from a (3+1)D topological quantum field theory (TQFT): the (3+1)D $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ gauge theory  [^5] The action of this TQFT in the continuum is $S_{\mathbb{Z}_{2}^{3}}=\pi\int_{\mathcal{M}^{4}}a^{(1)}db^{(1)}+a^{(2)}db^{(2)}+a^{(3)}db^{(3)}.$ (20) Here, $a^{(i)}$ and $b^{(i)}$ represent the $\mathbb{Z}_{2}$ 1-form electric gauge fields and 2-form magnetic gauge fields defined on the triangulated (3+1)D space-time manifold $\mathcal{M}^{4}=\mathcal{M}^{3}\times S^{1}_{t}$ ($\mathcal{M}^{3}$ being the space manifold, and the circle $S^{1}_{t}$ being the time part). equivalent to a 3D color code $CC(\mathcal{L}_{c})$ or three identical copies of 3D $\mathbb{Z}_{2}$ toric codes $\operatorname*{\text{\raisebox{1.07639pt}{{$\bigotimes$}}}}_{i=1}^{3}TC(\mathcal{L})$ supported on an arbitrary triangulation $\mathcal{L}$ of an arbitrary 3-manifold [32]. We define the operator valued 1-cochain $a^{(i)}\in C^{1}(\mathcal{L};\mathbb{Z}_{2})$ with its eigenvalues belonging to $\{0,1\}$ and $i$ denotes the $i^{\text{th}}$ copy of the toric code. The coefficient of each edge (1-cell) $e$ in the 1-cochain corresponds to a Pauli-$Z$ operator via
@@
Note that the Pauli operator has eigenvalues $\pm 1$ instead. Physically, $a^{(i)}$ corresponds to the $\mathbb{Z}_{2}$ electric gauge field. Similarly, there also exists a 1-cochain $b^{(i)}$ corresponding to the magnetic $\mathbb{Z}_{2}$ gauge field, which corresponds to a Pauli-X operator on each edge $e$ as
@@ III.2 Logical CCZ as a 0-form symmetry
For a triangulated manifold, one can define the cup product $\cup$’ between a $p$-cochain $\alpha^{p}$ and $q$-cochain $\beta^{q}$ evaluated on a $(p+q)$-simplex $v_{0}v_{1}\cdots v_{p+q}]$ as [67]
@@
The cup-product evaluated on each 3-simplex $v_{0}v_{1}v_{2}v_{3}]$ ($v_{i}$ represents the four vertices in the simplex)
@@
which explicitly shows the corresponding constant-depth circuit composed of many CCZ gates coupling the qubits in three copies of toric codes located on edge $v_{0}v_{1}],[v_{1}v_{2}]$ and $v_{2}v_{3}]$ in each simplex $v_{0}v_{1}v_{2}v_{3}]$, as illustrated in Fig. 7.
Figure 7: Illustration of the evaluation of the cup product in Eq. (III.2). A CCZ gate is coupling the qubits located on red, blue and green edges, i.e., $v_{0}v_{1}],[v_{1}v_{2}],[v_{2}v_{3}]$ in each simplex labeled by $v_{0}v_{1}v_{2}v_{3}]$. The CCZ gates in neighboring simplices may have overlap and they hence compose a constant-depth local circuit instead of a strictly transversal gate.
@@
We consider an $X$-stabilizer $S^{(1)}_{X;v}$$\prod_{e\supset v}X^{(1)}(e)$ located at vertex $v$ in the 1st copy of the toric code. We then consider conjugating $U$ with $S^{(1)}_{X;v}$:
@@
The condition in Eq. (37) should be considered as the gauge-invariance condition in the code space, since $a^{(1)}\rightarrow a^{(1)}+d\tilde{v}$ is just a gauge transformation in the $\mathbb{Z}_{2}$ gauge theory  [^6] This is in analogy to the famous gauge transformation $\vec{a}\rightarrow\vec{a}+\nabla\chi$ in Maxwell’s theory or the corresponding gauge theory in the continuum, where $\vec{a}$ is the quantum vector potential. The gauge invariance of the magnetic field $B=\nabla\times\vec{a}=0$ is satisfied due to the Stokes theorem: $\nabla\times\nabla\chi=0$.
@@
$P_{\mathcal{C}}\int_{{}_{\mathcal{L}}}d\tilde{v}\cup a^{(2)}P_{\mathcal{C}}=$ $P_{\mathcal{C}}\left[\int_{{}_{\mathcal{L}}}d(\tilde{v}\cup a^{(2)})+\int_{{}_{\mathcal{L}}}\tilde{v}\cup da^{(2)}\right]P_{\mathcal{C}}$
@@
Mathematically, the gauge invariance condition is associated with the topological invariance of $U$ with respect to arbitrary deformation of the cocycles by coboundaries.
@@
$a^{(1)}=\sum_{\alpha^{1}}\hat{n}_{\alpha}\alpha^{1},\quad a^{(2)}=\sum_{\beta^{1}}\hat{m}_{\beta}\beta^{1},\quad a^{(3)}=\sum_{\gamma^{1}}\hat{l}_{\gamma}\gamma^{1},$ (41)
where the quantum variable $\hat{n}_{\alpha}$, $\hat{m}_{\beta}$ and $\hat{l}_{\gamma}$ with eigenvalues $\{0,1\}$ are the winding numbers, i.e., coefficients of the cocycle basis. We can hence re-express $U$ as
$U=$ $\prod_{\alpha^{1},\beta^{1},\gamma^{1}}(-1)^{\int_{\mathcal{L}}(\hat{n}_{\alpha}\alpha^{1})\cup(\hat{m}_{\beta}\beta^{1})\cup(\hat{l}_{\gamma}\gamma^{1})}$ (42)
$=$ $\prod_{\alpha^{1},\beta^{1},\gamma^{1}}\left[(-1)^{\hat{n}_{\alpha}\hat{m}_{\beta}\hat{l}_{\gamma}}\right]^{\int_{\mathcal{L}}\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}$ (43)
$=$ $\prod_{\alpha^{1},\beta^{1},\gamma^{1}}\overline{\text{CCZ}}[(\alpha^{1};1),(\beta^{1};2),(\gamma^{1};3)]^{\int_{\mathcal{L}}{\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}}$ (44)
$=$ $\prod_{\alpha_{2},\beta_{2},\gamma_{2}}\overline{\text{CCZ}}[(\alpha_{2};1),(\beta_{2};2),(\gamma_{2};3)]^{{|\alpha_{2}\cup\beta_{2}\cup\gamma_{2}|}},$ (45)
where the third equality has use the relation $(-1)^{\hat{n}_{\alpha}\hat{m}_{\beta}\hat{l}_{\gamma}}$$\overline{\text{CCZ}}$. Here, $(\alpha^{1};1)$, $(\beta^{1};2)$ and $(\gamma^{1};3)$ are the labels of the logical qubit using the cocycle basis and the copy number. The logical CCZ gate between logical qubits $(\alpha^{1};1)$, $(\beta^{1};2)$ and $(\gamma^{1};3)$ is only non-trivial (not logical identity) if and only if $\int_{\mathcal{L}}{\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}$ evaluates non-trivially, i.e., have a non-trivial triple intersection of the Poincare dual: $|\alpha_{2}\cup\beta_{2}\cup\gamma_{2}|=1$.